# Solar carbon abundance from 3D non-LTE modelling of the diagnostic lines of the CH molecule

### v1

@@ 2.6 Spectrum Synthesis
We compute spectra for three different carbon abundances spaced by $0.4$ dex centred around $8.43$. Non-LTE spectra share a single set of departure coefficients, determined at the central abundance. For synthesis of spectra at intermediate abundances we use monotonic Hermite interpolation between the three reference spectra. We tested the impact of this interpolation method in 1D NLTE against computing departure coefficients and spectra directly with a given carbon abundance and found that it does not exceed $0.01$ dex in the abundance range between $8.35$ and $8.83$ dex for both optical and infrared lines (see Fig. 15 in the Appendix). ColletAsplund2007 showed that, in contrast to more metal-poor stars, the formation of CH is independent of the oxygen abundance in the Sun.
@@ 3 Results
Table 1: A-X line data employed in the CLV analysis. Most of the 9 discussed diagnostic transitions have multiple components. $E_{i}$ represents the lower excitation potential of each transition. $W_{\rm eq}$ represents the equivalent widths measured for each line in the KPNO solar flux atlas.
@@ 3.1.1 Continuum normalisation
Therefore we have decided to determine a linear correction to the continuum of the IAG center-to-limb observations in the wavelength region of $1.5\text{\,}\mathrm{\text{Å}}$ around the theoretical line centers using our 3D-NLTE synthetic spectra. For this we model the surrounding atomic lines in LTE using line data extracted from VALD (KupkaPiskunov1999; PiskunovKupka1995; RyabchikovaPiskunov2015). Problematic spectral lines, which are not well reproduced by the models, are masked out via sigma-clipping, considerably improving the continuum fit. We allow the carbon abundance to be varied in this initial fitting phase, but we only keep the continuum parameters. Strictly speaking the continuum correction is model dependent, but we find that the continuum levels predicted by the M3DIS and STAGGER models are indistinguishable. Having obtained the continuum correction we renormalise the continuum and determine the equivalent widths of the diagnostic lines by fitting a Gaussian line profile to the wavelength region of $0.08\text{\,}\mathrm{\text{Å}}$ around the theoretical line centers using SciPy’s curve_fit function. We take great care to estimate realistic uncertainties on the resulting equivalent widths measurements. The dominant error in the equivalent widths measurements of all lines is the continuum normalisation. The details of this error analysis are discussed in A.3.
@@ 3.1.2 Abundance fitting and center-to-limb variation
Having measured equivalent widths and their respective uncertainties for each IAG observation of a specific line, we fit the equivalent widths of the synthetic line profiles for all angles simultaneously. The employed line data for the A-X transitions is tabulated in Table 1 and the resulting fits are presented in Fig. 6, which illustrates the results obtained for spatially resolved IAG solar spectra. For comparison, we also show the results obtained using the KPNO solar disc-center intensity (brault1987spectral) and flux atlas (KuruczFurenlid1984) in Appendix 23. The typical error on the abundance due to the fitting routine is $\pm 0.04$ dex and it is approximately the same for all models as it is dominated by the uncertainty in the continuum placement. We note that overall the scatter of EW measurements is dominated by the properties of the observed IAG data (see also Fig. 16 in Appendix). In particular, the errors bars are larger at angles closer to the limb, which is due to the reduced brightness at the limb. Also for selected angles closer to the centre, such as that at $\mu=0.97,0.98$ ($r/r_{\odot}=0.24,0.20$), the observations are less reliable, because the data rely on a few observations only. This is, in particular, visible in the bottom left panel of Fig. 6 that shows the 4274 Å line. Clearly, the uncertainties of EW measurements are of the order $\sim 5$ percent, whereas the amplitude of the spatially-resolved data all CH lines spans a much larger range of over 20 percent and more importantly, the behaviour of the CH EWs is systematic, with EWs steadily increasing towards the limb. Thus, the quality of the IAG data and the behaviour of the EWs as a function of angle is sufficient to enable detailed diagnostics of our line formation models for CH and of C abundances, as we demonstrate below.
@@ 4 Comparison with previous studies
Our line selection overlaps with the one by AGS05 who focused on 9 A-X transitions and 102 infrared rotation-vibration transitions to determine the C abundance based on 3D LTE synthesis of CH lines. After the CH data by MasseronPlez2014 became available this analysis was updated for 7 out of the 9 A-X and 51 out of the 102 infrared transitions in AsplundAmarsi2021 (AAG21). Besides the updated line data AAG21 also employed the partition functions by BarklemCollet2016 and STAGGER snapshots with improved microphysics (MagicCollet2013) compared to AGS05. Both studies supplied abundances based on a 1D MARCS model for reference. An independent study by AlexeevaMashonkina2015 (AM15) also derived the solar carbon abundance employing MARCS models and investigating the same 9 A-X transitions as AGS05.
@@ 4.2 Infrared rotation-vibration transitions
AAG21 state a 3D LTE abundance of $8.47\pm 0.04$ as derived from 51 of the 102 rotation-vibration transitions employed in AGS05. Unfortunately there is no mention of how the original sample from AGS05 was cut in half. As the line-to-line scatter in AGS05 covers a range of roughly $0.2$ dex (Fig 4. in (AsplundGrevesse2005)) a systematic reduction of the sample can easily account for the difference between their result and our 3D LTE abundance of $8.51\pm 0.07$ derived with a solar STAGGER simulation similar to theirs. Additionally, the continuum level in the raw Spacelab-3 ATMOS data frequently exceeds unity (see Fig. 17 in the Appendix), which contributes another source of significant uncertainty as custom renormalisation is inevitable.
@@ 4.3 Granulation effects on CH number densities
As shown in Fig 7 we find only small differences in the EWs of the CH lines between the full 3D M3DIS model and its optically averaged <3D> version when removing all velocity components from the line formation calculation. This means that inhomogeneities in the full 3D model have a negligible effect on the CH line strength, at least in the Sun. This is an initially surprising result as UitenbroekCriscuoli2011 conclude that <3D> models always underestimate molecular number densities present in stellar atmospheres. They showcase this employing the CO molecule and find confirmation in the studies by KiselmanNordlund1995 and ScottAsplund2006, who derive larger abundances from <3D> models compared to full 3D models for OH and CO lines, respectively.
To explore this interesting behaviour further, we compare the difference between the <3D> number densities $N_{\rm<3D>}$ and the full 3D number densities averaged over surfaces of equal optical depth $N_{\rm 3D}$> for different molecules using a single M3DIS 3D RHD snapshot (see Fig. 20 in the Appendix). Using the <3D> (time and spatially averaged) atmosphere results in underestimating the number densities in the line forming regions for the CO, OH, NH and O2 molecules. However, in the case of CH and C2 the number densities are overestimated at $\log\tau<-0.5$. The reason for this is the high dissociation energy of the CO molecule, which is roughly three times as large as the dissociation energy of CH. This causes a substantial fraction of the total number of C atoms $N_{\rm C}$ to be locked in CO in this optical depth range. Below a certain temperature, the fraction of C locked in CH no longer increases with decreasing temperature, but is reduced due to $N_{\rm CO}/N_{\rm C}$ approaching unity. Thus, the number density of C i and therefore CH becomes anti-correlated to the number density of CO and the approximation – atomic number densities are constant under small perturbations in temperatures – used in UitenbroekCriscuoli2011 breaks down.
@@ 4.4 Solar Carbon abundance
The most recent analysis of atomic C i lines in the solar photosphere was conducted by AmarsiBarklem2019 (hereafter AB19), using 3D NLTE radiative transfer models similar to those employed in this work. AB19 determined a C abundance of $8.44\pm 0.02$, which is notably lower than our measurement of $8.52\pm 0.07$, despite their small uncertainties. However, AB19 noted that adopting the line selection and equivalent width measurements from CaffauLudwig2010 would increase their derived C abundance by $0.07$ dex. CaffauLudwig2010 performed a 3D LTE analysis of the solar C i diagnostics using CO5BOLD models. Furthermore, updated theoretical calculations of transition data for C ions by LiAmarsi2021 yielded a C abundance of $8.50\pm 0.07$ when the oscillator strengths from their study were applied to the AB19 analysis. AB19 state that 1D to 3D abundance differences (3D typically yielding higher abundances) are caused by differences in the atmospheric stratification finding that 3D-⟨3D⟩ abundance corrections are much less severe (<0.03 dex at disk-centre). MaggBergemann2022 find a C abundance of $8.56\pm 0.05$ for C i lines assuming LTE and employing an <3D> STAGGER model. This result agrees with our findings, especially when applying the 3D NLTE - 3D LTE abundance difference of 0.01-0.01 dex found by AB19. In sum, we show that our 3D NLTE C abundance of $8.52\pm 0.07$ based on molecular CH lines is in agreement with the C i measurements.
@@ 5 Discussion
Our new 3D NLTE solar C abundance is highly relevant within the context of Standard Solar Models (SSMs). These models are used, among other applications, to predict the rates of nuclear fusion processes in the Sun’s core (e.g. VinyolesSerenelli2017). The Sun generates energy through two primary mechanisms: the proton-proton (pp) chain, which is responsible for approximately 99% of the energy, and the much subdominant carbon-nitrogen-oxygen (CNO) cycle. Both processes produce neutrinos that easily escape the core, providing a direct probe of the physics of the Sun’s interior. Importantly, the CNO cycle, catalyzed by the abundances of C and N in the core, yields distinct neutrino fluxes (specifically from the decay of 13N, 15O, and 17F, Fig. 10) that are nearly linearly dependent on the sum of abundances of C and N in the core (e.g. Serenelli2013; AppelBagdasarian2022; BasilicoBellini2023). These neutrino fluxes, along with those produced in the pp chain can be probed in experiments, such as Borexino (BorexinoCollaborationAltenmuller2020; AppelBagdasarian2022). To explore how the properties of the solar core change with our new 3D NLTE A(C), we use our well-established and carefully-tested SSM (Bahcall2006; SerenelliBasu2009; SerenelliHaxton2011; VinyolesSerenelli2017) to compute theoretical neutrino fluxes for the Sun. For other elements, we use our data from BergemannLodders2025 and LoddersBergemann2025 (hereafter, BLP25).
@@
Fig. 10 shows our predicted neutrino fluxes for 7Be, 8B, and for the total CNO cycle (13N+15O+17F). The plotted contours outline the 68% confidence levels of our SSM-based neutrino fluxes. Here we also overplot the experimental data from the global analysis by Gonzalez-GarciaMaltoni2024. As suggested by the authors, we use their results derived without imposing the solar luminosity constraint [^2] We note that imposing the solar luminosity constraint leads to an overall reduction of neutrino fluxes from the Sun. However, it effectively leads to an increased carbon abundance, due to an anti-correlation between the fluxes released by the proton-electron-proton (pep) reaction and by the CNO cycle., as these results provide a purely measurement-based reference and do not rely on the SSM. We also include, for visual comparison, our results for the neutrino fluxes obtained using the composition by AsplundAmarsi2021. As seen in this Fig., the experimental solar neutrino fluxes are consistent with the predictions of the SSM based on our solar 3D NLTE A(C) and BLP25 abundances. However, there is a systematic discrepancy with the SSM neutrino results based on AAG21 composition, which has also been demonstrated and extensively discussed in AppelBagdasarian2022 and BasilicoBellini2023. We note that BLP25 provides more conservative errors on individual abundance measurements, hence the contours cover the largest areas.
@@
We note that the neutrino fluxes produced in a given nuclear reaction are strongly correlated with the temperature in the solar core $T_{c}$, which is an output of the SSM, and with the abundances of the elements (BahcallUlmer1996). The black dashed lines in Fig. 10 depict how the fluxes change with the change of $T_{c}$. This is crucial, because the uncertainties of all other input parameters of the SSM including abundances, nuclear reaction rates, and environmental parameters ($L_{\odot}$, opacity, age, and diffusion) enter the predicted fluxes from 7Be and 8B only through $T_{c}$. The only exception being the reactions 7Be+p and 7Be+e. However, these nuclear rates are well-understood and the measurement uncertainties are within $5\%$ (AcharyaAliotta2025). Hence, the impressive agreement between our model fluxes and the experimental data for 7Be and 8B fluxes - to better than 10% - indicates that the solar $T_{c}$ is well reproduced. In the case of CNO fluxes, once $T_{c}$ is fixed by the good agreement of 7Be and 8B, only two remaining factors can reconcile the AAG21-based SSM model predictions with the CNO measurement: (i) a higher C+N abundance, or (ii) a change in the 14N+p reaction rate. However, since the CNO neutrino flux depends almost linearly on this rate, a $\sim 40\%$ increase of the rate would be necessary to match the observations. This large difference is well-beyond the uncertainty of this reaction rate of $\sim 8\%$ (AcharyaAliotta2025). For a detailed analysis of the SSM we refer the reader to our earlier papers (e.g. Serenelli2013; VinyolesSerenelli2017, and references therein).
@@
Figure 11: Combined carbon and nitrogen abundance measurements from different components of the Sun. We adopted the nitrogen abundances of $7.94\pm 0.11$ (3D NLTE, BergemannLodders2025) and $7.88\pm 0.12$ (1D LTE, MaggBergemann2022) as the respective photospheric values. Here the ’Neutrino flux’ data correspond to the independent solar surface C+N value, obtained by combining the neutrino fluxes measured by the Borexino experiment (Gonzalez-GarciaMaltoni2024) and the SSM model (Serenelli2013).

### v2

@@ 2.6 Spectrum Synthesis
We compute spectra for three different carbon abundances spaced by $0.4$ dex centred around $8.43$. Non-LTE spectra share a single set of departure coefficients, determined at the central abundance. For synthesis of spectra at intermediate abundances we use monotonic Hermite interpolation between the three reference spectra. We tested the impact of this interpolation method in 1D NLTE against computing departure coefficients and spectra directly with a given carbon abundance and found that it does not exceed $0.01$ dex in the abundance range between $8.35$ and $8.83$ dex for both optical and infrared lines (see Fig. 15 in the Appendix). Collet et al. (2007) showed that, in contrast to more metal-poor stars, the formation of CH is independent of the oxygen abundance in the Sun.
@@ 3 Results
Table 1: A-X line data employed in the CLV analysis. Most of the 9 discussed diagnostic transitions have multiple components. $E_{i}$ represents the lower excitation potential of each transition. $W_{\rm eq}$ represents the equivalent widths measured for each feature (multiple blended components) in the KPNO solar flux atlas.
@@ 3.1.1 Continuum normalisation
Therefore we have decided to determine a linear correction to the continuum of the IAG center-to-limb observations in the wavelength region of $1.5\text{\,}\mathrm{\text{Å}}$ around the theoretical line centers using our 3D-NLTE synthetic spectra. For this we model the surrounding atomic lines in LTE using line data extracted from VALD (Kupka et al., 1999; Piskunov et al., 1995; Ryabchikova et al., 2015). Problematic spectral lines, which are not well reproduced by the models, are masked out via sigma-clipping, considerably improving the continuum fit. We allow the carbon abundance to be varied in this initial fitting phase, but we only keep the continuum parameters. Strictly speaking the continuum correction is model dependent, but we find that the continuum levels predicted by the M3DIS and STAGGER models are indistinguishable. Having obtained the continuum correction we renormalise the continuum and determine the equivalent widths of the diagnostic lines by fitting a Gaussian line profile to the wavelength region of $0.08\text{\,}\mathrm{\text{Å}}$ around the theoretical line centers using SciPy’s curve_fit function. We take great care to estimate realistic uncertainties on the resulting equivalent widths measurements. The dominant error in the equivalent widths measurements of all lines is the continuum normalisation. The details of this error analysis are discussed in A.3.
@@ 3.1.2 Abundance fitting and center-to-limb variation
Having measured equivalent widths and their respective uncertainties for each IAG observation of a specific line, we fit the equivalent widths of the synthetic line profiles for all angles simultaneously. The employed line data for the A-X transitions is tabulated in Table 1 and the resulting fits are presented in Fig. 6, which illustrates the results obtained for spatially resolved IAG solar spectra. For comparison, we also show the results obtained using the KPNO solar disc-center intensity (Brault and Neckel, 1987) and flux atlas (Kurucz et al., 1984) in Appendix 23. The typical error on the abundance due to the fitting routine is $\pm 0.04$ dex and it is approximately the same for all models as it is dominated by the uncertainty in the continuum placement. We note that overall the scatter of EW measurements is dominated by the properties of the observed IAG data (see also Fig. 16 in Appendix). In particular, the errors bars are larger at angles closer to the limb, which is due to the reduced brightness at the limb. Also for selected angles closer to the centre, such as that at $\mu=0.97,0.98$ ($r/r_{\odot}=0.24,0.20$), the observations are less reliable, because the data rely on a few observations only. This is, in particular, visible in the bottom left panel of Fig. 6 that shows the 4274 Å line. Clearly, the uncertainties of EW measurements are of the order $\sim 5$ percent, whereas the amplitude of the spatially-resolved data all CH lines spans a much larger range of over 20 percent and more importantly, the behaviour of the CH EWs is systematic, with EWs steadily increasing towards the limb. Thus, the quality of the IAG data and the behaviour of the EWs as a function of angle is sufficient to enable detailed diagnostics of our line formation models for CH and of C abundances, as we demonstrate below.
@@ 4 Comparison with previous studies
Our line selection overlaps with the one by AGS05 who focused on 9 A-X transitions and 102 infrared rotation-vibration transitions to determine the C abundance based on 3D LTE synthesis of CH lines. After the CH data by Masseron et al. (2014) became available this analysis was updated for 7 out of the 9 A-X and 51 out of the 102 infrared transitions in Asplund et al. (2021) (AAG21). Besides the updated line data AAG21 also employed the partition functions by Barklem and Collet (2016) and STAGGER snapshots with improved microphysics (Magic et al., 2013) compared to AGS05. Both studies supplied abundances based on a 1D MARCS model for reference. An independent study by Alexeeva and Mashonkina (2015) (AM15) also derived the solar carbon abundance employing MARCS models and investigating the same 9 A-X transitions as AGS05.
@@ 4.2 Infrared rotation-vibration transitions
AAG21 state a 3D LTE abundance of $8.47\pm 0.04$ as derived from 51 of the 102 rotation-vibration transitions employed in AGS05. Unfortunately there is no mention of how the original sample from AGS05 was cut in half. As the line-to-line scatter in AGS05 covers a range of roughly $0.2$ dex (Fig 4. in (Asplund et al., 2005)) a systematic reduction of the sample can easily account for the difference between their result and our 3D LTE abundance of $8.51\pm 0.07$ derived with a solar STAGGER simulation similar to theirs. Additionally, the continuum level in the raw Spacelab-3 ATMOS data frequently exceeds unity (see Fig. 17 in the Appendix), which contributes another source of significant uncertainty as custom renormalisation is inevitable.
@@ 4.3 Granulation effects on CH number densities
As shown in Fig 7 we find only small differences in the EWs of the CH lines between the full 3D M3DIS model and its optically averaged <3D> version when removing all velocity components from the line formation calculation. This means that inhomogeneities in the full 3D model have a negligible effect on the CH line strength, at least in the Sun. This is an initially surprising result as Uitenbroek and Criscuoli (2011) conclude that <3D> models always underestimate molecular number densities present in stellar atmospheres. They showcase this employing the CO molecule and find confirmation in the studies by Kiselman and Nordlund (1995) and Scott et al. (2006), who derive larger abundances from <3D> models compared to full 3D models for OH and CO lines, respectively.
To explore this interesting behaviour further, we compare the difference between the <3D> number densities $N_{\rm<3D>}$ and the full 3D number densities averaged over surfaces of equal optical depth $N_{\rm 3D}$> for different molecules using a single M3DIS 3D RHD snapshot (see Fig. 20 in the Appendix). Using the <3D> (time and spatially averaged) atmosphere results in underestimating the number densities in the line forming regions for the CO, OH, NH and O2 molecules. However, in the case of CH and C2 the number densities are overestimated at $\log\tau<-0.5$. The reason for this is the high dissociation energy of the CO molecule, which is roughly three times as large as the dissociation energy of CH. This causes a substantial fraction of the total number of C atoms $N_{\rm C}$ to be locked in CO in this optical depth range. Below a certain temperature, the fraction of C locked in CH no longer increases with decreasing temperature, but is reduced due to $N_{\rm CO}/N_{\rm C}$ approaching unity. Thus, the number density of C i and therefore CH becomes anti-correlated to the number density of CO and the approximation – atomic number densities are constant under small perturbations in temperatures – used in Uitenbroek and Criscuoli (2011) breaks down.
@@ 4.4 Solar Carbon abundance
The most recent analysis of atomic C i lines in the solar photosphere was conducted by Amarsi et al. (2019) (hereafter AB19), using 3D NLTE radiative transfer models similar to those employed in this work. AB19 determined a C abundance of $8.44\pm 0.02$, which is notably lower than our measurement of $8.52\pm 0.07$, despite their small uncertainties. However, AB19 noted that adopting the line selection and equivalent width measurements from Caffau et al. (2010) would increase their derived C abundance by $0.07$ dex. Caffau et al. (2010) performed a 3D LTE analysis of the solar C i diagnostics using CO5BOLD models. Furthermore, updated theoretical calculations of transition data for C ions by Li et al. (2021) yielded a C abundance of $8.50\pm 0.07$ when the oscillator strengths from their study were applied to the AB19 analysis. AB19 state that 1D to 3D abundance differences (3D typically yielding higher abundances) are caused by differences in the atmospheric stratification finding that 3D-⟨3D⟩ abundance corrections are much less severe (<0.03 dex at disk-centre). Magg et al. (2022) find a C abundance of $8.56\pm 0.05$ for C i lines assuming LTE and employing an <3D> STAGGER model. This result agrees with our findings, especially when applying the 3D NLTE - 3D LTE abundance difference of 0.01-0.01 dex found by AB19. In sum, we show that our 3D NLTE C abundance of $8.52\pm 0.07$ based on molecular CH lines is in agreement with the C i measurements.
@@ 5 Discussion
Our new 3D NLTE solar C abundance is highly relevant within the context of Standard Solar Models (SSMs). These models are used, among other applications, to predict the rates of nuclear fusion processes in the Sun’s core (e.g. Vinyoles et al., 2017). The Sun generates energy through two primary mechanisms: the proton-proton (pp) chain, which is responsible for approximately 99% of the energy, and the much subdominant carbon-nitrogen-oxygen (CNO) cycle. Both processes produce neutrinos that easily escape the core, providing a direct probe of the physics of the Sun’s interior. Importantly, the CNO cycle, catalyzed by the abundances of C and N in the core, yields distinct neutrino fluxes (specifically from the decay of 13N, 15O, and 17F, Fig. 10) that are nearly linearly dependent on the sum of abundances of C and N in the core (e.g. Serenelli et al., 2013; Appel et al., 2022; Basilico et al., 2023). These neutrino fluxes, along with those produced in the pp chain can be probed in experiments, such as Borexino (Borexino Collaboration et al., 2020; Appel et al., 2022). To explore how the properties of the solar core change with our new 3D NLTE A(C), we use our well-established and carefully-tested SSM (Bahcall et al., 2006; Serenelli et al., 2009, 2011; Vinyoles et al., 2017) to compute theoretical neutrino fluxes for the Sun. For other elements, we use our data from Bergemann et al. (2025) and Lodders et al. (2025) (hereafter, BLP25).
@@
Fig. 10 shows our predicted neutrino fluxes for 7Be, 8B, and for the total CNO cycle (13N+15O+17F). The plotted contours outline the 68% confidence levels of our SSM-based neutrino fluxes. Here we also overplot the experimental data from the global analysis by Gonzalez-Garcia et al. (2024). As suggested by the authors, we use their results derived without imposing the solar luminosity constraint [^2] We note that imposing the solar luminosity constraint leads to an overall reduction of neutrino fluxes from the Sun. However, it effectively leads to an increased carbon abundance, due to an anti-correlation between the fluxes released by the proton-electron-proton (pep) reaction and by the CNO cycle., as these results provide a purely measurement-based reference and do not rely on the SSM. We also include, for visual comparison, our results for the neutrino fluxes obtained using the composition by Asplund et al. (2021). As seen in this Fig., the experimental solar neutrino fluxes are consistent with the predictions of the SSM based on our solar 3D NLTE A(C) and BLP25 abundances. However, there is a systematic discrepancy with the SSM neutrino results based on AAG21 composition, which has also been demonstrated and extensively discussed in Appel et al. (2022) and Basilico et al. (2023). We note that BLP25 provides more conservative errors on individual abundance measurements, hence the contours cover the largest areas.
@@
We note that the neutrino fluxes produced in a given nuclear reaction are strongly correlated with the temperature in the solar core $T_{c}$, which is an output of the SSM, and with the abundances of the elements (Bahcall and Ulmer, 1996). The black dashed lines in Fig. 10 depict how the fluxes change with the change of $T_{c}$. This is crucial, because the uncertainties of all other input parameters of the SSM including abundances, nuclear reaction rates, and environmental parameters ($L_{\odot}$, opacity, age, and diffusion) enter the predicted fluxes from 7Be and 8B only through $T_{c}$. The only exception being the reactions 7Be+p and 7Be+e. However, these nuclear rates are well-understood and the measurement uncertainties are within $5\%$ (Acharya et al., 2025). Hence, the impressive agreement between our model fluxes and the experimental data for 7Be and 8B fluxes - to better than 10% - indicates that the solar $T_{c}$ is well reproduced. In the case of CNO fluxes, once $T_{c}$ is fixed by the good agreement of 7Be and 8B, only two remaining factors can reconcile the AAG21-based SSM model predictions with the CNO measurement: (i) a higher C+N abundance, or (ii) a change in the 14N+p reaction rate. However, since the CNO neutrino flux depends almost linearly on this rate, a $\sim 40\%$ increase of the rate would be necessary to match the observations. This large difference is well-beyond the uncertainty of this reaction rate of $\sim 8\%$ (Acharya et al., 2025). For a detailed analysis of the SSM we refer the reader to our earlier papers (e.g. Serenelli et al., 2013; Vinyoles et al., 2017, and references therein).
@@
Figure 11: Combined carbon and nitrogen abundance measurements from different components of the Sun. We adopted the nitrogen abundances of $7.94\pm 0.11$ (3D NLTE, Bergemann et al., 2025) and $7.88\pm 0.12$ (1D LTE, Magg et al., 2022) as the respective photospheric values. Here the ’Neutrino flux’ data correspond to the independent solar surface C+N value, obtained by combining the neutrino fluxes measured by the Borexino experiment (Gonzalez-Garcia et al., 2024) and the SSM model (Serenelli et al., 2013).